40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 162     Entry time: Fri Aug 30 12:22:56 2013
Author: Koji 
Type: Optics 
Category: General 
Subject: H1 OMC Cavity side UV gluing 

H1 OMC Cavity side optics was glued on the breadboard

Curved mirror gluing

- Applied the UV glues to CM1/CM2 prisms.

- Checked the spot positions on the curved mirrors

- Apply 50V to CM1

- Measure the FSR and TMS while the cavity was locked.

FSR: 261.70925MHz
TMS_V: 57.60500MHz
TMS_H: 57.94125MHz

=> Cavity round trip length of 1.1455m
=> TMS/FSR = {0.220111, 0.221395}

First accidental resonance is the lower sideband at 28th order modes.

Carrier 9th-order HOM: 2.9~7.6 line width away
Upper Sideband 13th-order HOM: 14.1-20.7 LW away
Lower Sideband 19th-order HOM: 3.3-13.1 LW away

- As this result was satisfactory, the UV illumination was zapped. It did not change the alignment. The cavity was kept locked during the illumination.

Peripheral optics gluing

- QPD path BS/Steering Mirrors were glued
- DCPD path BS was glued

The UV glue was applied to the optics.
Then the optics were placed on the breadboard along with the fixture.

Placed the dummy QPD/DCPD mount with the alignment disks.
The horizontal positions of the spots were well with in the horizontal range of the mounts.
 The UV illumination was zapped. Checked the alignment again and no problem was found.

ELOG V3.1.3-